# Developed and Quasi-Developed Macro-Scale Heat Transfer in Micro- and Mini-Channels with Arrays of Offset Strip Fins Subject to a Uniform Heat Flux

### v1

@@ Abstract
Micro-and Mini-Channels, Offset Strip Fin Array, Macro-Scale Modeling, Quasi-Developed Heat Transfer, Closure
@@ 1 Introduction
Our previous studies [14, 15] highlighted that the flow regime in micro- and mini-channels with arrays of offset strip fins is mainly laminar and steady [5, 6, 7, 8, 9, 10, 11, 12, 13, 16]. The corresponding heat transfer regime in these channels is characterized by the flow’s Prandtl number, which is usually 0.7 or 7 since air and water are the most common working fluids [5, 6, 7, 8, 9, 12, 13, 16]. Nevertheless, the Prandtl number of cryogenic fluids like helium, hydrogen, and nitrogen, both in the gaseous and liquid state, is of the same order of magnitude as that of air [10, 11]. Furthermore, the heat transfer regime in micro- and mini-channel applications is commonly investigated under the assumption of a steady and uniform heat flux at the channel wall. This boundary condition is considered to represent well the heat transfer process in microelectronics cooling systems, balanced counter-flow heat recuperators, as well as solar air heaters, and heat exchangers in cryogenic systems [11, 12, 13, 17, 18, 19, 20, 21, 22].
@@
On the contrary, no numerical or analytical studies exist in the literature on the development of heat transfer in channels with arrays of fins or other solid structures. Further progress on the topic appears to be hampered by the tremendously high computational cost of DNS for resolving the flow and temperature fields in large arrays [25]. As a result, the general features of the developing flow and heat transfer in such channels are still barely understood, with the exception of the quasi-periodically developed flow regime [54]. The former regime is characterized by a single exponential velocity mode with a streamwise periodic amplitude and dominates over almost the entire entrance region in micro- and mini-channels with offset strip fins [40] and inline cylinders [55]. Although the existence of a similar quasi-periodically developed heat transfer regime has also been recently deduced [56], its significance within the entire region of developing heat transfer has yet to be investigated, as empirical data is still lacking.
@@
Not only theoretical work, but also experimental work on the heat transfer development in channels with arrays of solid structures remains scarce. To the authors’ knowledge, the study by Dong et al. [57] provides the only experimental investigation of the influence of thermal development on the thermo-hydraulic performance of offset strip fin arrays. Dong et al. present a correlation for the Colburn j-factor, which represents the overall heat transfer coefficient in the channel measured through temperature probes upstream and downstream of various air-side offset strip fin arrays. In this correlation, the influence of the developing flow and heat transfer regime has been taken into account via a scaling factor based on the ratio of the total channel length to the offset strip fin length. It should be noted, though, that the correlation mainly covers data for air flows in the transitional regime and applies only to conventional offset strip fin channels with relatively high channel heights. Moreover, as the heat transfer data was obtained using high flow rates of hot water on the control side, the correlation reflects a constant wall temperature condition rather than a constant heat flux boundary condition. Therefore, the analysis of Dong et al. is of limited significance for micro- and mini-channels with offset strip fin arrays, particularly when the heat flux at the wall is imposed.
@@
From the preceding literature survey we conclude that, currently, the importance of the heat transfer development in micro- and mini-channels with offset strip fin arrays is still unexplored. Therefore, this work numerically assesses to what extent the heat transfer regime can be described as developed on a macro-scale level in common micro- and mini-channels with offset strip fins subject to a uniform heat flux. Specifically, we analyze the onset point of the developed heat transfer regime, as well as the onset point of quasi-developed heat transfer inside these channels. Hereto, we rely on DNS to resolve the entire temperature field in the simultaneously developing channel flow for Reynolds numbers ranging from 28 to 1224. Both air and water are considered as working fluids, each in combination with a highly thermally conductive solid material, such as copper. Further, we assess the accuracy of the Nusselt number correlations from our preceding work [15] in the region where the heat transfer regime is still developing, since they are strictly speaking only valid for predicting the heat transfer coefficient in the developed region. Concretely, we conduct a comparison with the actual macro-scale heat transfer coefficient along the channel, which we compute from the full temperature field through a discretized double volume-averaging operator. Finally, we investigate and quantify the influence of the channel’s side walls on the developed macro-scale temperature field and the actual macro-scale heat transfer coefficient.
@@ 2.1 Geometry of the channel and fin array
In this work, the following array sizes have been selected for our DNS simulations: $N_{1}=20$, $N_{2}\in\{5,10,13,15,17\}$. Furthermore, an inlet and outlet region without fins over a length $s_{0}=l_{1}$ and $s_{N}=2.5l_{1}$ have been added in front and behind the offset strip fin array. These parameter values are the same as those considered in our previous study on the macro-scale flow development in offset strip fin arrays [58]. They have been chosen such that our results are representative of typical geometries of micro- and mini-channels with offset strip fins [5, 6, 7, 8, 9, 10, 11, 12, 13]. We have verified that the number of fin rows $N_{1}=20$ is large enough to accurately assess the development lengths for both the flow and temperature fields. Besides, we have checked that the number of lateral unit cells $N_{2}$ results in true lateral periodicity of the flow field at a distance $l_{2}$ from the channel’s sidewalls.
@@ 3.1 Onset of periodically developed heat transfer
Accordingly, we define the onset point $x_{\text{periodic,T}}$ of the periodically developed heat transfer regime as the streamwise coordinate $x_{1}$ after which the temperature field agrees with the periodically developed solution (8) within 10% of the overall temperature change over a single unit cell, so $|T-T_{\text{dev}}|/|\mathrm{\nabla{T}}\cdot\bm{l}_{1}|\leqslant 0.1$ for $x_{1}\geqslant x_{\text{periodic,T}}$. The practical advantage of this definition lies in the fact the value of $x_{\text{periodic,T}}$ remains independent of the total array length. Moreover, due to the employed criterion for defining $x_{\text{periodic,T}}$, we can generally expect expression (8) to hold within at least 1% of the total temperature difference over the entire developed region. The explanation is that the developed region reasonably covers more than 10 unit cells in the streamwise direction, given the high number of fin rows in contemporary micro- and mini-channels [5, 6, 7, 8, 9, 10, 11, 12, 13]. Another interesting consequence of the chosen criterion is illustrated later in this section: For all the cases examined in this study, the local macro-scale heat transfer coefficient and local Nusselt number along the centerline of the channel are observed to deviate at most 5% from their constant values in the developed region, when this region is identified as $x_{1}\geqslant x_{\text{periodic,T}}$.
@@
Figure 4 shows that the relation between the onset point $x_{\text{periodic,T}}$ and the Prandtl number $Pr_{f}$ can be reduced to a linear dependence on the Péclet number $Pe\triangleq Re_{b}Pr_{f}$. This appears to be true for all geometries and flow conditions studied in this work. We found, for instance, that the data in Figure 4 is captured by the following linear correlations to within a maximum error of 5%: $(x_{\text{periodic,T}}-s_{0})/l_{1}\simeq 0.0214Pe-0.899$ when $t/l=0.02$, and $(x_{\text{periodic,T}}-s_{0})/l_{1}\simeq 0.0198Pe-0.680$ when $t/l=0.04$. It should be noted though that the former linear correlations are only valid for $Pe>42$ and $Pe>34$, respectively. For smaller values of the Péclet number, the onset point approximately coincides with the start of the fin array: $x_{\text{periodic,T}}\simeq s_{0}$, as it can be seen in Figure 4.
@@ 3.2 Region of developed macro-scale heat transfer
According to Figure 7(b), this observation appears to be quite generally valid over a substantial range of Reynolds numbers. In particular, all our DNS results indicate that the deviations from the developed temperature profiles remain lower than 20% of the temperature difference over a single unit cell $|\mathrm{\nabla{T}}\cdot\bm{l}_{1}|$ in the region behind the first row of fins, up to the last row of the fins. On that account, we conclude that, from a practical perspective, the region of developed macro-scale heat transfer coincides with the region over which the weighted porosity $\epsilon_{fm}$ is independent of the streamwise coordinate $x_{1}$, i.e. $x_{1}\in\left(x_{\text{in}},x_{\text{out}}\right)$ with $x_{\text{in}}=s_{0}+l_{1}$ and $x_{\text{out}}=L_{1}-\left(s_{N}+l_{1}\right)$. Outside the former region, hence immediately near the channel inlet and outlet, the temperature field does become affected by the specific boundary conditions, as well as the presence of strong porosity gradients.
Figure 7: Intrinsic fluid and solid macro-scale temperature outside (black) and inside (grey) the side-wall region, when $Pr_{f}=7$, $Re_{b}=28$, $h/l=0.12$, $s/l=0.48$, $t/l=0.02$, $s_{0}=l_{1}$, $s_{N}=2.5l_{1}$, $N_{1}=20$, and $N_{2}=10$ (a), and intrinsic fluid macro-scale temperature along the channel centerline ($x_{2}=L_{2}/2$), when $Pr_{f}=7$, $h/l=0.12$, $s/l=0.48$, $t/l=0.02$, $s_{0}=l_{1}$, $s_{N}=2.5l_{1}$, $N_{1}=20$, and $N_{2}=10$ (b)
@@ 3.3 Accuracy of the developed Nusselt number correlation
In principle, the criterion $h_{b}\simeq h_{\text{unit}}$ will result in a different value for the onset point of the developed regime (11) in comparison with the former criterion (9). However, according to our numerical results, both criteria are practically identical, as Figure 8 reveals.
@@
In Figure 8, we show the actual macro-scale heat transfer coefficient $h_{b}(\bm{x})$ along a channel at four different lateral positions with an array of $20\times 10$ offset strip fin unit cells. In this figure, we have represented $h_{b}(\bm{x})$ by the local Nusselt number $Nu_{\text{b}}(\bm{x})\triangleq h_{b}l^{2}/k_{f}$. The two subfigures at the top correspond to an air flow, while the two subfigures at the bottom correspond to a water flow. The considered Reynolds numbers for both fluids are relatively low ($Re_{b}=42$ and $Re_{b}=56$), so that the resulting Péclet numbers are low for air ($Pe\simeq 40$) and moderate for water ($Pe\simeq 400$), as of interest for many micro- and mini-channel applications [70, 7, 8].
@@
Figure 8 illustrates that the actual macro-scale heat transfer coefficient $h_{b}$ deviates less than 5% from its constant value for $x_{1}\geq x_{\text{periodic,T}}=x_{\text{dev,T}}-l_{1}$, when viewed along the centerline $x_{2}=5l_{2}$ of the channel. This demonstrates that the onset point $x_{\text{periodic,T}}$ is also a good measure to indicate where the developed Nusselt number correlation becomes valid, even though $x_{\text{periodic,T}}$ is based on a criterion for the underlying temperature field instead (8). Due to numerical errors, however, the constant value of $h_{b}$ in the developed value region still differs slightly from the constant value $h_{\text{unit}}$ predicted by the correlations (12) and (13): the difference $|h_{b}-h_{\text{unit}}|/h_{b}$ may be up to 7%. These numerical errors originate mainly from the limited accuracy of the correlation, as the discretization error on $h_{b}$ is around 1%.
@@
Figure 8 also gives an impression of the typical accuracy of the developed Nusselt number correlations in the region where the flow and temperature fields are developing, at least for low to moderate Péclet numbers. In the figure, we have compared the actual local macro-scale heat transfer coefficient $h_{b}(\bm{x})$ with its local prediction $h_{\text{unit}}(\bm{x})$ based on the local Nusselt number $Nu_{\text{unit}}\left(\bm{x}\right)$. Hereto, we have evaluated $Nu_{\text{unit}}\left(\bm{x}\right)$ based on the local Reynolds number $Re_{l}\left(\bm{x}\right)\triangleq\rho_{f}\|\langle\bm{u}\rangle_{m}\|l/\mu_{f}$, for the given Prandtl number and unit cell geometry. Essentially, this means that we treat the macro-scale heat transfer regime locally as almost developed, so that the local macro-scale heat transfer coefficient is governed by the periodically developed temperature equations from [34], even when the flow is still developing. The main takeaway from this figure is that the local value $h_{\text{unit}}\left(\bm{x}\right)$ predicted by the developed Nusselt number correlation is also quite accurate over a significant part of the developing flow and heat transfer region. After all, even upstream of the strict onset point $x_{\text{dev,T}}$, the developed Nusselt number correlations (12)-(13) predict the actual macro-scale heat transfer coefficient $h_{b}$ with a mean and maximum relative error $|h_{b}-h_{\text{unit}}|/h_{b}$ of 10% and 25%, respectively. One part of the explanation is that the actual macro-scale heat transfer coefficient in the developing flow and heat transfer region only deviates between 5% and 30% from its constant value in the developed region. However, the small error between the actual heat transfer coefficient $h_{b}(\bm{x})$ and its local approximation $h_{\text{unit}}\left(\bm{x}\right)$ immediately upstream of the point $x_{\text{dev,T}}$ can also be explained in a different way. As we will show in the next section, this error is determined by the inherent characteristics of the quasi-periodically developed heat transfer regime in $\Omega_{\text{predev,T}}$.
@@
Although Figure 8 applies to a single channel geometry and four Péclet numbers, we expect that these results are indicative for a wide range of applications [5, 6, 7, 8, 9, 10, 11, 12, 13] in which the channel height is low and the Reynolds numbers remain moderate. So, we expect that our developed heat transfer correlations can adequately model the macro-scale heat transfer regime in micro- and mini-channels with offset strip fins for most engineering purposes. Certainly, for the conditions studied here, the effects of developing heat transfer are negligible at the macro-scale level, when an accuracy level of 10% to 20% for the heat transfer coefficient is deemed acceptable. In that case, the modeling errors due to the use of developed Nusselt number correlations are also larger than the actual deviations from the developed heat transfer coefficient in the developing region anyway.
@@
Figure 8: Macro-scale heat transfer coefficient $Nu_{b}$ (full) and its prediction $Nu_{\text{unit}}$ (dashed) by the developed correlations from [15], when $h/l=0.12$, $s/l=0.48$, $t/l=0.02$, $s_{0}=l_{1}$, $s_{N}=2.5l_{1}$, $N_{1}=20$, and $N_{2}=10$

### v2

@@ Abstract
Micro-and Mini-Channels, Offset Strip Fin Array, Macro-Scale Modelling, Quasi-Developed Heat Transfer, Closure
@@ 1 Introduction
Our previous studies [14, 15] highlighted that the flow regime in micro- and mini-channels with arrays of offset strip fins is mainly laminar and steady [5, 6, 7, 8, 9, 10, 11, 12, 13, 16]. The corresponding heat transfer regime in these channels is characterized by the flow’s Prandtl number, which is usually 0.7 or 7 since air and water are the most common working fluids. Nevertheless, the Prandtl number of cryogenic fluids like helium, hydrogen, and nitrogen, both in the gaseous and liquid state, is of the same order of magnitude as that of air [10, 11]. Furthermore, the heat transfer regime in micro- and mini-channel applications is commonly investigated under the assumption of a steady and uniform heat flux at the channel wall. This boundary condition is considered to represent well the heat transfer process in microelectronics cooling systems, balanced counter-flow heat recuperators, as well as solar air heaters, and heat exchangers in cryogenic systems [11, 12, 13, 17, 18, 19, 20, 21, 22].
@@
On the contrary, no analytical or numerical studies exist in the literature on the development of heat transfer in channels with arrays of fins or other solid structures. It remains an open problem whether analytical methods for such channels can be formulated, as the available methods cannot easily be extended to deal with the more complex boundary conditions in three-dimensional arrays. The research gap in numerical studies is primarily due to the tremendously high computational cost of simulating the developing flow and temperature fields in large arrays [25]. Even for steady laminar flows, which do not require turbulence modelling via Reynolds-Averaged Navier–Stokes (RANS) simulations or Large Eddy Simulations (LES) and allow the use of DNS, the main computational challenge remains the required mesh resolution, hence the global mesh size. In typical fin arrays, a minimal mesh size of the order of $10^{5}$ cells per unit cell is required for a sufficient spatial resolution of the flow and heat transfer [54, 14, 15]. This makes DNS, as well as RANS or LES in case the developing flow would be turbulent, computationally prohibitive for such fin arrays. As a result, the general features of the developing flow and heat transfer in such channels are still barely understood, with the exception of the quasi-periodically developed flow regime [55]. The former regime is characterized by a single exponential velocity mode with a streamwise periodic amplitude and dominates over almost the entire entrance region in micro- and mini-channels with offset strip fins [40] and inline cylinders [56]. Although the existence of a similar quasi-periodically developed heat transfer regime has also been recently deduced [57], its significance within the entire region of developing heat transfer has yet to be investigated, as empirical data is still lacking.
@@
Not only theoretical work, but also experimental work on the heat transfer development in channels with arrays of solid structures remains scarce. The most significant study, to the authors’ knowledge, is by Dong et al. [58], who investigated the impact of thermal development on the thermo-hydraulic performance of offset strip fin arrays. They proposed a correlation for the Colburn j-factor, accounting for developing flow and heat transfer through a scaling factor based on the channel length-to-fin length ratio. However, this correlation primarily applies to air flows in the transitional regime and is specifically tailored to conventional offset strip fin channels with relatively high channel heights. Moreover, since the study involved high flow rates of hot water on the control side, it reflects a constant-wall-temperature condition. As a result, the findings of Dong et al. have limited applicability to our investigation of heat transfer in micro- and mini-channels with offset strip fins under a constant heat flux boundary condition.
@@
From the preceding literature survey we conclude that, currently, the importance of the heat transfer development in micro- and mini-channels with offset strip fin arrays is still unexplored. Therefore, the objective of this work is to theoretically and numerically assess to what extent the heat transfer regime can be described as developed on a macro-scale level in common micro- and mini-channels with offset strip fins subject to a uniform heat flux. Specifically, we analyze the onset point of the developed heat transfer regime, as well as the onset point of quasi-developed heat transfer inside these channels. Hereto, we rely on DNS to resolve the entire temperature field in the simultaneously developing channel flow for Reynolds numbers ranging from 28 to 1224. Both air and water are considered as working fluids, each in combination with a highly thermally conductive solid material, such as copper. Further, we assess the accuracy of the Nusselt number correlations from our preceding work [15] in the region where the heat transfer regime is still developing, since they are strictly speaking only valid for predicting the heat transfer coefficient in the developed region. Concretely, we conduct a comparison with the actual macro-scale heat transfer coefficient along the channel, which we compute from the full temperature field through a discretized double volume-averaging operator. Finally, we investigate and quantify the influence of the channel’s side walls on the developed macro-scale temperature field and the actual macro-scale heat transfer coefficient.
@@ 2.1 Geometry of the channel and fin array
In this work, the following array sizes have been selected for our DNS simulations: $N_{1}=20$, $N_{2}\in\{5,10,13,15,17\}$. Furthermore, an inlet and outlet region without fins over a length $s_{0}=l_{1}$ and $s_{N}=2.5l_{1}$ have been added in front and behind the offset strip fin array. These parameter values are the same as those considered in our previous study on the macro-scale flow development in offset strip fin arrays [59]. They have been chosen such that our results are representative of typical geometries of micro- and mini-channels with offset strip fins [59]. We have verified that the number of fin rows $N_{1}=20$ is large enough to accurately assess the development lengths for both the flow and temperature fields. Besides, we have checked that the number of lateral unit cells $N_{2}$ results in true lateral periodicity of the flow field at a distance $l_{2}$ from the channel’s sidewalls.
@@ 3.1 Onset of periodically developed heat transfer
Accordingly, we define the onset point $x_{\text{periodic,T}}$ of the periodically developed heat transfer regime as the streamwise coordinate $x_{1}$ after which the temperature field agrees with the periodically developed solution (8) within 10% of the overall temperature change over a single unit cell, so $|T-T_{\text{dev}}|/|\mathrm{\nabla{T}}\cdot\bm{l}_{1}|\leqslant 0.1$ for $x_{1}\geqslant x_{\text{periodic,T}}$. The practical advantage of this definition lies in the fact the value of $x_{\text{periodic,T}}$ remains independent of the total array length. Moreover, due to the employed criterion for defining $x_{\text{periodic,T}}$, we can generally expect expression (8) to hold within at least 1% of the total temperature difference over the entire developed region. The explanation is that the developed region reasonably covers more than 10 unit cells in the streamwise direction, given the high number of fin rows in contemporary micro- and mini-channels [59]. Another interesting consequence of the chosen criterion is illustrated later in this section: For all the cases examined in this study, the local macro-scale heat transfer coefficient and local Nusselt number along the centerline of the channel are observed to deviate at most 5% from their constant values in the developed region, when this region is identified as $x_{1}\geqslant x_{\text{periodic,T}}$.
@@
Figure 4 shows that the relation between the onset point $x_{\text{periodic,T}}$ and the Prandtl number $Pr_{f}$ can be reduced to a linear dependence on the Péclet number $Pe\triangleq Re_{b}Pr_{f}$. This appears to be true for all geometries and flow conditions studied in this work. We found, for instance, that the data in Figure 4 is captured by the following linear correlations within a maximum error of 5%: $(x_{\text{periodic,T}}-s_{0})/l_{1}\simeq 0.0214Pe-0.899$ when $t/l=0.02$, and $(x_{\text{periodic,T}}-s_{0})/l_{1}\simeq 0.0198Pe-0.680$ when $t/l=0.04$. It should be noted though that the former linear correlations are only valid for $Pe>42$ and $Pe>34$, respectively. For smaller values of the Péclet number, the onset point approximately coincides with the start of the fin array: $x_{\text{periodic,T}}\simeq s_{0}$, as it can be seen in Figure 4.
@@ 3.2 Region of developed macro-scale heat transfer
According to Figure 7(b) and Figure 7(c), this observation appears to be quite generally valid over a substantial range of Reynolds numbers. In particular, all our DNS results indicate that the deviations from the developed temperature profiles remain lower than 20% of the temperature difference over a single unit cell $|\mathrm{\nabla{T}}\cdot\bm{l}_{1}|$ in the region behind the first row of fins, up to the last row of the fins. On that account, we conclude that, from a practical perspective, the region of developed macro-scale heat transfer coincides with the region over which the weighted porosity $\epsilon_{fm}$ is independent of the streamwise coordinate $x_{1}$, i.e. $x_{1}\in\left(x_{\text{in}},x_{\text{out}}\right)$ with $x_{\text{in}}=s_{0}+l_{1}$ and $x_{\text{out}}=L_{1}-\left(s_{N}+l_{1}\right)$. Outside the former region, hence immediately near the channel inlet and outlet, the temperature field does become affected by the specific boundary conditions, as well as the presence of strong porosity gradients.
Figure 7: Intrinsic fluid and solid macro-scale temperature outside (black) and inside (grey) the side-wall region, when $Pr_{f}=7$, $Re_{b}=28$, $h/l=0.12$, $s/l=0.48$, $t/l=0.02$, $s_{0}=l_{1}$, $s_{N}=2.5l_{1}$, $N_{1}=20$, and $N_{2}=10$ (a). Intrinsic fluid macro-scale temperature (b) and its gradient (c) along the channel centerline ($x_{2}=L_{2}/2$), when $Pr_{f}=7$, $h/l=0.12$, $s/l=0.48$, $t/l=0.02$, $s_{0}=l_{1}$, $s_{N}=2.5l_{1}$, $N_{1}=20$, and $N_{2}=10$. Note that the non-dimensional gradient along the main flow direction is noted by $\nabla^{+}_{1}\triangleq\bm{l}_{1}\cdot\nabla$
@@ 3.3 Accuracy of the developed Nusselt number correlation
In principle, the criterion $h_{b}\simeq h_{\text{unit}}$ will result in a different value for the onset point of the developed regime (11) in comparison with the former criterion (9). However, according to our numerical results, both criteria are practically identical, as Figures 8 and 9 reveal.
@@
In Figures 8 and 9, we show the actual macro-scale heat transfer coefficient $h_{b}(\bm{x})$ along a channel at four different lateral positions with an array of $20\times 10$ offset strip fin unit cells. In this figure, we have represented $h_{b}(\bm{x})$ by the local Nusselt number $Nu_{\text{b}}(\bm{x})\triangleq h_{b}l^{2}/k_{f}$. Figure 8 corresponds to an air flow, while Figure 9 corresponds to a water flow. The considered Reynolds numbers for both fluids are relatively low to moderate ($Re_{b}=28-112$), so that the resulting Péclet numbers are low to moderate for air ($Pe\simeq 30-80$) and moderate for water ($Pe\simeq 200-400$), as of interest for many micro- and mini-channel applications [71, 7, 8].
@@
Figures 8 and 9 illustrate that the actual macro-scale heat transfer coefficient $h_{b}$ deviates less than 5% from its constant value for $x_{1}\geq x_{\text{periodic,T}}=x_{\text{dev,T}}-l_{1}$, when viewed along the centerline $x_{2}=5l_{2}$ of the channel. This demonstrates that the onset point $x_{\text{periodic,T}}$ is also a good measure to indicate where the developed Nusselt number correlation becomes valid, even though $x_{\text{periodic,T}}$ is based on a criterion for the underlying temperature field instead (8). Due to numerical errors, however, the constant value of $h_{b}$ in the developed value region still differs slightly from the constant value $h_{\text{unit}}$ predicted by the correlations (12) and (13): the difference $|h_{b}-h_{\text{unit}}|/h_{b}$ may be up to 7%. These numerical errors originate mainly from the limited accuracy of the correlation, as the discretization error on $h_{b}$ is around 1%.
@@
Figures 8 and 9 also give an impression of the typical accuracy of the developed Nusselt number correlations in the region where the flow and temperature fields are developing, at least for Péclet numbers in relevant micro- and mini-channel applications. In the figure, we have compared the actual local macro-scale heat transfer coefficient $h_{b}(\bm{x})$ with its local prediction $h_{\text{unit}}(\bm{x})$ based on the local Nusselt number $Nu_{\text{unit}}\left(\bm{x}\right)$. Hereto, we have evaluated $Nu_{\text{unit}}\left(\bm{x}\right)$ based on the local Reynolds number $Re_{l}\left(\bm{x}\right)\triangleq\rho_{f}\|\langle\bm{u}\rangle_{m}\|l/\mu_{f}$, for the given Prandtl number and unit cell geometry. Essentially, this means that we treat the macro-scale heat transfer regime locally as almost developed, so that the local macro-scale heat transfer coefficient is governed by the periodically developed temperature equations from [34], even when the flow is still developing. The main takeaway from this figure is that the local value $h_{\text{unit}}\left(\bm{x}\right)$ predicted by the developed Nusselt number correlation is also quite accurate over a significant part of the developing flow and heat transfer region. After all, even upstream of the strict onset point $x_{\text{dev,T}}$, the developed Nusselt number correlations (12)-(13) predict the actual macro-scale heat transfer coefficient $h_{b}$ with a mean and maximum relative error $|h_{b}-h_{\text{unit}}|/h_{b}$ of 10% and 25%, respectively. One part of the explanation is that the actual macro-scale heat transfer coefficient in the developing flow and heat transfer region only deviates between 5% and 30% from its constant value in the developed region. However, the small error between the actual heat transfer coefficient $h_{b}(\bm{x})$ and its local approximation $h_{\text{unit}}\left(\bm{x}\right)$ immediately upstream of the point $x_{\text{dev,T}}$ can also be explained in a different way. As we will show in the next section, this error is determined by the inherent characteristics of the quasi-periodically developed heat transfer regime in $\Omega_{\text{predev,T}}$.
Given the large range of Péclet numbers covered, we expect that the former results are indicative of a wide range of practical applications, especially those where the channel height-to-length ratio $h/l$ is below 0.2 and the Reynolds number remains below 150 [15]. After all, for such geometries, the Péclet number has the most significant effect on the onset of developed heat transfer, as shown in Section 3.1.
The most severe limitation of the results in Figures 8 and 9 is the assumed parabolic inlet velocity profile for the flow simulations [59]. For strongly asymmetric inlet flow, the former results will no longer be representative. However, in practical applications, channel manifolds are designed to establish ideal operating conditions, effectively ensuring a more uniform inlet velocity profile. Such more uniform profiles induce smaller velocity perturbations near the channel inlet than numerically predicted, according to the experimental measurements of Eneren et al. [72]. In turn, these smaller velocity perturbations imply that also the deviations from the developed heat transfer regime will be even smaller in practice, across different channel geometries. Together with the broad Péclet number range considered, this reinforces the generality of our conclusions, supporting the applicability of our findings across various micro- and mini-channel configurations with offset strip fins.
@@
For the former reasons, we thus expect that our developed heat transfer correlations can adequately model the macro-scale heat transfer regime in the most commonly applied micro- and mini-channels with offset strip fins. Certainly, for the conditions studied here, the effects of developing heat transfer are negligible at the macro-scale level when an accuracy level of 10% to 20% for the heat transfer coefficient is deemed acceptable. In that case, the modelling errors due to the use of developed Nusselt number correlations are larger than the actual deviations from the developed heat transfer coefficient in the developing region.
@@
Figure 8: Macro-scale heat transfer coefficient $Nu_{b}$ (full) and its prediction $Nu_{\text{unit}}$ (dashed) by the developed correlations from [15], when $Pr_{f}=0.7$, $h/l=0.12$, $s/l=0.48$, $t/l=0.02$, $s_{0}=l_{1}$, $s_{N}=2.5l_{1}$, $N_{1}=20$, and $N_{2}=10$
Figure 9: Macro-scale heat transfer coefficient $Nu_{b}$ (full) and its prediction $Nu_{\text{unit}}$ (dashed) by the developed correlations from [15], when $Pr_{f}=7$, $h/l=0.12$, $s/l=0.48$, $t/l=0.02$, $s_{0}=l_{1}$, $s_{N}=2.5l_{1}$, $N_{1}=20$, and $N_{2}=10$